OpenStudy (anonymous):

Examine the function f(x) = (xe^[(.5)(x-2)^2])*(x+2)^3 Use logarithmic differentiation to calculate the derivative of f(x) . At what values does logarithmic differentiation fail?

OpenStudy (anonymous):

\[f(x)=(xe^{(.5)(x-2)^2})*(x+2)^3 \] \[ln(f(x))=ln((xe^{(.5)(x-2)^2})*(x+2)^3)=ln(x) + ln(e^{(.5)(x-2)^2}) + ln((x+2)^3)\] \[=ln(x) + (.5)(x-2)^2 + 3*ln(x+2)\] Apply the rules of logarithms to this function. \[ln(\frac{x}{y})=ln(x) - ln(y)\] \[ln(x*y)=ln(x) + ln(y)\] \[ln(x^n)=n*ln(x)\] \[ln(e^(x^n))=x^n\] You should be able to differentiate more easily now.

OpenStudy (anonymous):

OH! Okay, I was going to say this is ridiculous.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

What would be the best method of finding the tangent line of this ugly thing?

OpenStudy (anonymous):

A tangent line is basically in the form of y=mx + b to put it in better variables, t(a) = f'(a)*x + (f(a)-f'(a)*a) Where your y-intercept (b=f(a)-f'(a)*a)

OpenStudy (anonymous):

right but finding where f '(x) = 0 would be terrible